OpenStudy (anonymous):

Miguel's insurance company will replace his car if repair costs exceed 80% of the car's value. The car recently sustained $11600 in damage, but it was not replaced. What was the value of his car? a) >= $145000? b) $14500? c) More than 9280? or d) Less than $14500? I thought it was "c" but that doesn't seem right.

OpenStudy (saifoo.khan):

i have done this before! :(

OpenStudy (anonymous):

\[ \$11600 \le 0.8x, x \ge \$14500 \]

OpenStudy (anonymous):

less than 14500 80%=11600 100%=11600/80 *100=14500 as mentionned, car insurance is not replaced, its value must be less than 14500

OpenStudy (dumbcow):

car is greater than 14,500. It can't be less, if car is worth say 14,000, then 80% is .8*14,000 = 11,200 11,600 > 11,200, thus the car would have been replaced

hero (hero):

Slight oversight by john

hero (hero):

I lost track of you

OpenStudy (anonymous):

Yea, it took me longer than I thought and then I keep losing my internet. I am almost done though.